Preface

The Proceedings of the International Congress of Mathematicians 1998, held in Berlin, are published – electronically and in print – in three volumes. Volume I contains information on the organization of the Congress including the list of participants, reports on the opening and closing ceremonies, the Laudationes on the Fields Medalists and the Nevanlinna Prize Winner, and the Plenary Lectures. Volumes II and III contain the Invited Lectures.

For the first time, the Proceedings of an ICM have been produced completely electronically – without any commercial assistance. Using the facilities of Documenta Mathematica, the contents of the Plenary and Invited Lectures were made available without charge on the Internet, already before the Congress started, at http://www.mathematik.uni-bielefeld.de/documenta/.

The printed versions of Volumes II and III were distributed to the participants at the beginning of the Congress. Volume I, containing material which had to be gathered during the Congress, was printed about three months after the Congress.

We want to thank all the speakers and organizers for their cooperation which made such fast publication possible.

October 1998 Gerd Fischer Ulf Rehmann

Past Congresses

1897 Zürich 1900 Paris 1904 Heidelberg 1908 Roma 1912 Cambridge, UK 1920 Strasbourg 1924 Toronto 1928 Bologna 1932 Zürich 1936 Oslo 1950 Cambridge, USA 1954 Amsterdam 1958 Edinburgh 1962 Stockholm 1966 Moskva 1970 Nice 1974 Vancouver 1978 Helsinki 1982 Warszawa (held in 1983) 1986 Berkeley 1990 Kyoto 1994 Zürich

1998 Berlin

Past Fields Medalists and Rolf Nevanlinna Prize Winners

Recipients of Fields Medals

1936 Lars V. Ahlfors

Jesse Douglas

1950 Laurent Schwartz

Atle Selberg

1954 Kunihiko Kodaira

Jean-Pierre Serre

1958 Klaus F. Roth

René Thom

1962 Lars Hörmander

John W. Milnor

1966 Michael F. Atiyah

Paul J. Cohen Alexander Grothendieck Steve Smale

1970 Alan Baker

Heisuke Hironaka Sergei P. Novikov John G. Thompson

1974 Enrico Bombieri

David B. Mumford

1978 Pierre R. Deligne

Charles F. Fefferman Grigorii A. Margulis Daniel G. Quillen

1982 Alain Connes

William P. Thurston Shing-Tung Yau

1986 Simon K. Donaldson

Gerd Faltings Michael H. Freedman

1990 Vladimir G. Drinfeld

Vaughan F. R. Jones Shigefumi Mori Edward Witten

1994 Jean Bourgain

Pierre-Louis Lions Jean-Christophe Yoccoz Efim Zelmanov

Rolf Nevanlinna Prize Winners

1982 Robert E. Tarjan
1986 Leslie G. Valiant

1990 Alexander A. Razborov
1994 Avi Wigderson

14

Documenta Mathematica · Extra Volume ICM 1998 · I